OpenStudy (anonymous):

The equation, y = 19(1.057)x , represents book sales in billions of dollars since 1990. x represents the number of years after 1990. How many billions of dollars of books were sold in 1990? Use x = 0 and round the answer to the nearest whole number.

OpenStudy (anonymous):

y = 19(1.057)^x, where x = 20 y = 19(1.057)^20 = 19(3.03) = 57.57 $58 billion